# Learning-Based Modeling of a Magnetically Steerable Soft Suction Device for Endoscopic Endonasal Interventions

### v1

@@ III Experimental Setup
Figure 3: Experimental setup for magnetic actuation and shape sensing. The system includes an FBG-embedded soft robot mounted on a linear actuator, positioned above electromagnetic coils at a distance of $d_{tip}$. A power supply and stepper driver control actuation, while real-time strain data is collected via the FBG interrogator and visualized on a workstation. Close-ups on the right show the FBG clamping, suction tube, 3D printed fixtures, and the soft robot with a tip magnet for steering.
(a) 3D Shape (0.2 Hz)
(b) XY View (0.2 Hz)
(c) XY View (0.4 Hz)
(d) XY View (0.6 Hz)
(e) XY View (0.8 Hz)
(f) XY View (1.0 Hz)
(g) Magnetic Field Profiles (0.2 Hz)
Figure 4: Top: Workspace of the soft robot tip at 100 mm above the coil surface under various magnetic field strengths and frequencies. Each curve represents a centerline shape reconstructed from FBG data. Bottom right: Time-varying magnetic field components ($b_{x}$, $b_{y}$) for the 0.2 Hz trials that generated the tip motions shown in (b).
@@
The experimental setup was designed to evaluate the deformation behavior of the device under controlled magnetic fields while capturing real-time shape data through embedded FBG sensors. The soft suction device was fabricated using biocompatible SIL 30 material (Carbon Inc., Redwood City, California, USA) via digital light synthesis (DLS) 3D printing, ensuring precise geometry and material consistency suitable for medical applications. SIL 30 is a silicone urethane elastomer with a nominal Shore A hardness of 35 (instant), as reported by the manufacturer [27]. The device measured 40 mm in length, with an OD of 4 mm and an inner channel diameter of 2 mm. A hollow cylindrical permanent magnet was securely affixed to the distal tip of the device to enable magnetic steering. The magnetization profile of this tip magnet was configured as shown in Fig. 3, optimized for effective actuation within rotating magnetic fields. Then, the fiber was threaded through the suction channel to enable continuous shape sensing along the length of the soft robot. The proximal end of the fiber was fixed at the base, while the distal end was anchored at the robot’s tip to ensure consistent alignment and accurate strain transfer during deformation. Owing to the fiber’s small OD (0.38 mm), its integration within the lumen imposed negligible mechanical constraint on the flexibility and natural deformation of the soft robot. The proximal end of the MCF was clamped onto a linear actuator, which allowed fine control over the vertical positioning of the soft robot relative to the electromagnetic coil system’s surface. Both the soft robot and the linear actuator were mounted onto a custom-designed aluminum frame to maintain structural stability and ensure consistent experimental conditions throughout data collection. Additionally, two 3D-printed supports were employed to secure both the base of the FBG fiber and the soft robot to the linear actuator, minimizing unwanted vibrations or displacement. For visual monitoring, two Logitech cameras were positioned orthogonally, providing multi-angle views of the robot’s deformation during actuation.
[…]
@@
To illustrate the magnetic field conditions applied during soft robot actuation, we visualized the in-plane rotating magnetic field components ($b_{x}$, $b_{y}$) for the 0.2 Hz excitation frequency across magnetic field magnitudes from 2 mT to 14 mT ($b_{z}=0$). These sinusoidal field profiles correspond directly to the workspace visualizations presented in Fig. 4b, where tip deformation was analyzed under identical field conditions. As shown in Fig. 4g, the blue and red curves represent the $b_{x}$ and $b_{y}$ components, respectively, generating a circular rotating field in the XY plane. Initially, the robot’s tip was positioned 100 mm above the coil surface. Upon completion of the test series at this height, the linear actuator was used to adjust the tip position to 90 mm above the table surface, and the full set of experiments was repeated under identical conditions to assess the influence of vertical distance on soft robot performance. These two heights were selected to represent clinically relevant working distances that fall within the safe operating range of the electromagnetic coil system, while ensuring measurable differences in magnetic field strength and resulting deformation. In total, 70 distinct actuation scenarios were conducted, covering all combinations of magnetic field strengths, rotation frequencies, and the two vertical distances. Throughout these tests, the strain data from all 26 FBG gratings were continuously recorded at a sampling rate of 30 Hz using a custom MATLAB script interfaced via TCP with the optical interrogator. Following data acquisition, the recorded 250 mm sensing length of the MCF was processed in post-experiment analysis. Only the final 40 mm segment, which corresponds precisely to the embedded length within the soft robot, was retained for shape reconstruction and modeling.
IV EVALUATIONS AND DISCUSSION
@@
This section presents a comprehensive evaluation of the learning-based approaches used for modeling the device. We compare the performance of Neural Network (NN) and Random Forest (RF) models in predicting Bezier control points from magnetic field parameters, analyze feature importance, and evaluate shape reconstruction accuracy.
IV-A Dataset Characteristics and Model Training
@@
IV-B Model Performance Comparison
Table II presents the detailed performance metrics of both models across all control points. The RF model outperforms the NN model, achieving an overall 28.37% reduction in RMSE. This improvement is particularly pronounced for the distal control points ($\mathbf{p}_{2}$ and $\mathbf{p}_{3}$), which are critical for accurate device tip positioning.
TABLE II: Performance comparison between NN and RF.
@@ IV-B Model Performance Comparison
RMSE R2 Max Err RMSE R2 Max Err
(mm) - (mm) (mm) - (mm)
@@
The model accuracy comparison by individual control points, as shown in Fig. 6a, reveals that while both models predict proximal control points ($\mathbf{p}_{0}$ and $\mathbf{p}_{1}$) with similar accuracy, the RF model outperforms the NN model for distal control points ($\mathbf{p}_{2}$ and $\mathbf{p}_{3}$). Quantitatively, the RF model reduces RMSE by 5.97% for $\mathbf{p}_{0}$, 5.47% for $\mathbf{p}_{1}$, 16.41% for $\mathbf{p}_{2}$, and an impressive 53.94% for $\mathbf{p}_{3}$ compared to the NN model. The error distribution analysis reveals that both models exhibit primarily zero-centered errors (Fig. 6c), with the RF model demonstrating tighter error distributions across all control points. The NN model struggles with $\mathbf{p}_{3}$ coordinates, exhibiting error exceeding 1.5 mm for ${p}_{3Y}$, while the RF model maintains max absolute errors below 1.11 mm.
@@
(a) Prediction accuracy for all Bezier control points. Blue dots: ground truth, red circles: NN predictions, green circles: RF predictions.
@@
(b) Comparing NN-predicted vs. actual Bezier control points across training, validation, testing, and full datasets.
@@
(c) Error histogram showing the distribution of prediction errors. The concentration of errors around zero indicates unbiased predictions.
Figure 6: Comparison of regression performance and prediction accuracy across learning models. (a) Prediction accuracy for all control points. (b) Detailed, zoomed-in correlation plots demonstrating goodness of fit after normalization by the robot length. (c) Signed error histograms highlighting the tighter residual spread achieved by the RF model.
IV-C Feature Importance Analysis
@@
To gain insights into the relationship between magnetic field parameters and device deformation, we analyzed the feature importance metrics derived from the RF model. Our results revealed distinct patterns of influence across different control points: The magnetic field components ($b_{x}$ and $b_{y}$) dominate the influence on most control point coordinates, with $b_{y}$ showing particularly high importance ($>$50%) for ${p}_{0Y}$, ${p}_{2Y}$, and ${p}_{3Y}$ coordinates. This aligns with the physical understanding of magnetic actuation, where the Y-component of the magnetic field primarily controls lateral movement of the device. The frequency parameter exhibits uniquely high importance (32.34%) for ${p}_{0Z}$, suggesting that actuation frequency plays a significant role in determining the base height of the device. The distance parameter shows consistent influence on $\mathbf{p}_{0}$ coordinates (23.81% for ${p}_{0X}$, 24.50% for ${p}_{0Z}$) and $\mathbf{p}_{3}$ coordinates (13.00% for ${p}_{3X}$, 13.17% for ${p}_{3Y}$, 15.79% for ${p}_{3Z}$), reflecting how the proximity to the electromagnetic coil system affects both the fixed base and the free tip of the device. This feature importance analysis provides valuable insights for control system design, indicating which magnetic field parameters should be prioritized for precise manipulation of specific device coordinates. For instance, fine control of the device tip position would benefit from careful modulation of both $b_{x}$ and $b_{y}$, while actuation frequency could be leveraged for height adjustment at the device base.
IV-D Bezier Curve Shape Reconstruction
Beyond individual control point accuracy, we evaluated the overall shape reconstruction performance by comparing the Bezier curves generated from predicted control points with ground truth curves. Figure 7 illustrates this comparison from both 3D and top-view perspectives.
@@ IV-D Bezier Curve Shape Reconstruction
Figure 7: Bezier curve shape prediction comparison between ground truth (blue) and RF model predictions (green) shown from the 3D perspective (left) and side view (right).
[…]
TABLE III: Comparison of learning models on the test set.
@@
V CONCLUSION

### v2

@@ 3 Experimental Setup
Figure 3: Experimental setup for magnetic actuation and shape sensing of soft robot. The system includes an FBG-embedded soft robot mounted on a linear actuator, positioned above electromagnetic coils at a distance of $d_{tip}$. A power supply and stepper driver control the linear actuation, while real-time strain data is collected via the FBG interrogator and visualized on a workstation. Close-ups on the right show the FBG clamping, suction tube, 3D printed fixtures, and the soft robot with a tip magnet for steering.
@@
The experimental setup was designed to evaluate the deformation behavior of the soft robot under controlled magnetic fields while capturing real-time shape data through embedded FBG sensors. The soft suction device was fabricated using biocompatible SIL 30 material (Carbon Inc., Redwood City, California, USA) via digital light synthesis (DLS) 3D printing, ensuring precise geometry and material consistency suitable for medical applications. SIL 30 is a silicone urethane elastomer with a nominal Shore A hardness of 35 (instant), as reported by the manufacturer [31]. The device measured 40 mm in length, with an OD of 4 mm and an inner channel diameter of 2 mm. A hollow cylindrical neodymium permanent magnet (Grade N50, 6.35 mm OD $\times$ 3.18 mm ID $\times$ 3.18 mm thickness, nickel-coated) was securely affixed to the distal tip of the device to enable magnetic steering. The magnet features through-thickness magnetization as shown in Fig. 3, optimized for effective actuation within rotating magnetic fields.
The FBG fiber was threaded through the suction channel to enable continuous shape sensing along the length of the soft robot. The proximal end of the fiber was fixed at the base, while the distal end was anchored at the robot’s tip to ensure consistent alignment and accurate strain transfer during deformation. Owing to the fiber’s small OD (0.38 mm), its integration within the lumen imposed negligible mechanical constraint on the flexibility and natural deformation of the soft robot. The proximal end of the MCF was clamped onto a linear actuator, which allowed fine control over the vertical positioning of the soft robot relative to the electromagnetic coil system’s surface. Both the soft robot and the linear actuator were mounted onto a custom-designed aluminum frame to maintain structural stability and ensure consistent experimental conditions throughout data collection. Additionally, two 3D-printed supports were employed to secure both the base of the FBG fiber and the soft robot to the linear actuator, minimizing unwanted vibrations or displacement. For visual monitoring, two Logitech cameras were positioned orthogonally, providing multi-angle views of the robot’s deformation during actuation.
Figure 4: Time-varying magnetic field components ($b_{x}$, $b_{y}$) for the 0.2 Hz trials corresponding to the trajectories in Fig. 5.
Magnetic actuation was achieved using an electromagnetic navigation system comprising eight high-power coils arranged beneath the workspace [32]. This system was engineered to deliver spatially controllable magnetic fields while maximizing workspace accessibility. The coil system was capable of generating magnetic fields up to 38 mT along the $x$ and $y$ axes, and up to 47 mT along the $z$ axis at a working height of 120 mm above the coil surface. During each experimental run, an in-plane rotating magnetic field was applied parallel to the $xy$ plane to induce controlled bending and rotation of the robot’s tip (Fig. 4). The applied magnetic field vector followed a sinusoidal form given by:
$\mathbf{b}(t)=A[-\sin(\omega t),\ \cos(\omega t),\ 0]^{\top}$ (9)
where $A$ is the magnetic field magnitude (ranging from 0 to 14 mT) and $\omega=2\pi f$ is the angular frequency corresponding to the rotation frequency $f$ (0.2, 0.4, 0.6, 0.8, and 1 Hz). This dynamic formulation generated a uniform circular magnetic field in the $xy$ plane, producing a time-varying torque on the embedded tip magnet that resulted in cyclic bending motion. As shown in Fig. 4, the blue and red curves represent the $b_{x}$ and $b_{y}$ components, respectively, generating a circular rotating field in the XY plane. To illustrate the magnetic field conditions applied during soft robot actuation, we visualized the in-plane rotating magnetic field components ($b_{x}$, $b_{y}$) for the 0.2 Hz excitation frequency across magnetic field magnitudes from 2 mT to 14 mT ($b_{z}=0$). These sinusoidal field profiles correspond directly to the workspace visualizations presented in Fig.5, where tip deformation was analyzed under identical field conditions.
Figure 5: 3D shape reconstruction (0.2 Hz). Each curve represents a centerline shape reconstructed from FBG data at different magnetic field strengths.
@@
A dynamic field was deliberately applied to capture the frequency dependent deformation characteristics of the soft robot, which are influenced by the viscoelastic behavior of the silicone material. As shown in Figs. 6, the same magnetic field magnitude led to larger rotation radii at higher frequencies, highlighting the presence of dynamic effects. The chosen frequency range (0.2–1 Hz) was selected to reflect typical motion speeds observed during endoscopic procedures, ensuring the evaluation remained within clinically relevant operational conditions. To record the data, the robot’s tip was initially positioned 100 mm above the electromagnetic coil table surface. Upon completion of the test series at this height, the linear actuator was used to adjust the tip position to 90 mm above the table surface, and the full set of experiments was repeated under identical conditions to assess the influence of vertical distance on soft robot performance. These two heights were selected to represent clinically relevant working distances that fall within the safe operating range of the electromagnetic coil system, while ensuring measurable differences in magnetic field strength and resulting deformation. In total, 70 distinct actuation scenarios were conducted, covering all combinations of magnetic field strengths, rotation frequencies, and the two vertical distances. Throughout these tests, the strain data from all 26 FBG gratings were continuously recorded at a sampling rate of 30 Hz using a custom MATLAB script interfaced via TCP with the optical interrogator. Following data acquisition, the recorded 250 mm sensing length of the MCF was processed in post-experiment analysis. Only the final 40 mm segment, which corresponds precisely to the embedded length within the soft robot, was retained for shape reconstruction and modeling.
(a) XY View (0.2 Hz)
(b) XY View (0.4 Hz)
(c) XY View (0.6 Hz)
(d) XY View (0.8 Hz)
(e) XY View (1.0 Hz)
Figure 6: XY-view trajectories at five actuation frequencies. Each panel shows tip paths reconstructed from FBG data under varying magnetic field strengths.
4 EVALUATIONS AND DISCUSSION
@@
This section presents a comprehensive evaluation of the learning based approaches used for modeling the soft suction device. We compare the performance of Neural Network (NN) and Random Forest (RF) models in predicting Bezier control points from magnetic field parameters, analyze feature importance, and evaluate shape reconstruction accuracy in the following sections.
4.1 Dataset Characteristics and Model Training
@@
4.2 Model Performance Comparison
Table 2 presents the detailed performance metrics of both models across all control points. The RF model outperforms the NN model, achieving an overall 28.37% reduction in RMSE. This improvement is particularly pronounced for the distal control points ($\mathbf{p}_{2}$ and $\mathbf{p}_{3}$), which are critical for accurate device tip positioning. The model accuracy comparison by individual control points, as shown in Figs. 8 and 9, reveals that while both models predict proximal control points ($\mathbf{p}_{0}$ and $\mathbf{p}_{1}$) with similar accuracy, the RF model outperforms the NN model for distal control points ($\mathbf{p}_{2}$ and $\mathbf{p}_{3}$). Quantitatively, the RF model reduces RMSE by 5.97% for $\mathbf{p}_{0}$, 5.47% for $\mathbf{p}_{1}$, 16.41% for $\mathbf{p}_{2}$, and an impressive 53.94% for $\mathbf{p}_{3}$ compared to the NN model. The error distribution analysis reveals that both models exhibit primarily zero-centered errors (Fig. 10), with the RF model demonstrating tighter error distributions across all control points. The NN model struggles with $\mathbf{p}_{3}$ coordinates, exhibiting error exceeding 1.5 mm for ${p}_{3Y}$, while the RF model maintains max absolute errors below 1.11 mm.
Table 2: Performance comparison between Neural Network (NN) and Random Forest (RF) models.
@@ 4.2 Model Performance Comparison
RMSE R2 Max Error RMSE R2 Max Error
($m$) - ($m$) ($m$) - ($m$)
@@
Figure 8: Prediction accuracy for all Bezier control points. Blue dots: ground truth, red circles: NN predictions, green circles: RF predictions.
@@
Figure 9: Comparison of NN-predicted and actual Bezier control points across training, validation, testing, and full datasets.
@@
Figure 10: Error histogram showing the distribution of prediction errors. The concentration of errors around zero indicates unbiased predictions.
4.3 Feature Importance Analysis
@@
To gain insights into the relationship between magnetic field parameters and device deformation, we analyzed the feature importance metrics derived from the RF model. Our results revealed distinct patterns of influence across different control points: The magnetic field components ($b_{x}$ and $b_{y}$) dominate the influence on most control point coordinates, with $b_{y}$ showing particularly high importance (>>50%) for ${p}_{0Y}$, ${p}_{2Y}$, and ${p}_{3Y}$ coordinates. This aligns with the physical understanding of magnetic actuation, where the Y-component of the magnetic field primarily controls lateral movement of the device. The frequency parameter exhibits uniquely high importance (32.34%) for ${p}_{0Z}$, suggesting that actuation frequency plays a significant role in determining the base height of the device. The distance parameter shows consistent influence on $\mathbf{p}_{0}$ coordinates (23.81% for ${p}_{0X}$, 24.50% for ${p}_{0Z}$) and $\mathbf{p}_{3}$ coordinates (13.00% for ${p}_{3X}$, 13.17% for ${p}_{3Y}$, 15.79% for ${p}_{3Z}$), reflecting how the proximity to the electromagnetic coil system affects both the fixed base and the free tip of the device. This feature importance analysis provides valuable insights for control system design, indicating which magnetic field parameters should be prioritized for precise manipulation of specific device coordinates. For instance, fine control of the device tip position would benefit from careful modulation of both $b_{x}$ and $b_{y}$, while actuation frequency could be leveraged for height adjustment at the device base.
4.4 Bezier Curve Shape Reconstruction
Beyond individual control point accuracy, we evaluated the overall shape reconstruction performance by comparing the Bezier curves generated from predicted control points with ground truth curves. Figures 11 and 12 illustrate this comparison from both 3D and top-view perspectives. The 3D visualization confirms that the RF model achieves excellent shape reconstruction across diverse deformation patterns. Predicted Bezier curves follow the ground-truth trajectories closely, preserving control-point placement and smooth interpolation. Quantitatively, the RF delivers a MAE of $0.064\;\text{mm}$ and a $95^{\text{th}}$-percentile error of $0.225\;\text{mm}$, both negligible relative to the device diameter (4 mm) and therefore suitable for safe navigation around delicate skull-base anatomy during endoscopic endonasal procedures.
@@ 4.4 Bezier Curve Shape Reconstruction
Figure 11: Bezier curve shape prediction (3D view). Ground truth (blue) vs. RF predictions (green).
Figure 12: Bezier curve shape prediction (side view). Ground truth (blue) vs. RF predictions (green).
Table 3: Comparison of learning models on the test set.
@@
The regression analysis of the NN model (Fig. 9) yielded correlation coefficients of $R_{\text{train}}=0.9169$, $R_{\text{val}}=0.8815$, and $R_{\text{test}}=0.8782$, indicating good generalization with limited over-fitting. Nevertheless, the RF model surpassed the NN across every metric. Using the test set, we formed per-sample total errors $\sum_{i=1}^{12}(y_{i}-\hat{y}_{i})^{2}$ for both models. Lilliefors tests detected slight departures from perfect normality for NN and RF residuals ($p=0.001$ for each). Levene’s test showed unequal variances ($p=4.96\times 10^{-21}$); consequently, Welch’s two-sample $t$-test was applied, revealing a highly significant reduction in error for RF compared with NN ($t=12.55$, $p=1.32\times 10^{-34}$). Mean ($\pm$ SD) total error fell from $0.283\pm 0.329$ for the NN to $0.131\pm 0.204$ for the RF (Table 3). These findings confirm that model selection materially influences residual error and reinforce our choice of RF for subsequent real-time deployment.
4.5 FBG Ground Truth Limitations and Accuracy
The accuracy validation methodology of this study warrants careful examination, particularly regarding the use of reconstructed shapes from FBG as ground truth for training the learning based models. While multi-core fiber Bragg grating sensors are well-established for shape sensing in medical robotics applications, their inherent limitations must be acknowledged when interpreting the reported accuracy metrics. The FBG system employed in this study is notably coarser than comparable studies, such as Dong et al. [33], who achieved 3.17 mm grating spacing and validated their FBG measurements against electromagnetic tracking, reporting 1.53 mm tip position accuracy. The piecewise constant curvature assumption between grating points introduces interpolation errors that accumulate along the fiber length, potentially resulting in tip position uncertainties exceeding 1 mm despite the sensor’s high strain sensitivity.
The reported sub-millimeter accuracy (0.064 mm RMSE) in this work specifically quantifies the Random Forest model’s ability to predict Bezier control points that, when used to reconstruct the shape, closely match the FBG-derived measurements. This metric should not be interpreted as absolute shape sensing accuracy, but rather as the model’s capacity to learn and reproduce the mapping between magnetic field inputs and FBG-measured deformations. The distinction is critical: while the learning-based approach successfully captures the nonlinear relationship between actuation and measured response with high fidelity, the absolute accuracy of the predicted shapes relative to true physical deformation remains bounded by the FBG system’s inherent limitations. Future work should incorporate independent validation using calibrated vision systems or higher resolution tracking methods to establish absolute accuracy bounds. Despite these limitations, the learning-based approach remains valid for the intended application of closed-loop control, where consistent relative shape prediction is more critical than absolute position accuracy. The model’s ability to predict deformation patterns with high repeatability enables effective navigation control even if absolute position uncertainty exists, as the controller can compensate for systematic errors through feedback mechanisms.
Furthermore, this study utilized only a single physical prototype with data collected across 70 actuation scenarios, preventing assessment of device-to-device variability or manufacturing tolerances. The learning-based model is inherently device-specific, trained on this particular prototype’s characteristics without validation across multiple devices. Clinical translation would necessitate either individual calibration for each manufactured device or training across multiple prototypes to establish model robustness. While the current work demonstrates the feasibility of the learning-based approach for shape prediction, comprehensive validation across multiple devices with varying material properties and manufacturing variations remains essential for establishing clinical applicability and generalization capabilities of the proposed methodology.
Lastly, the current study trained and evaluated the model exclusively using uniform, symmetric rotating magnetic fields generated under controlled laboratory conditions. Clinical environments present significantly more complex magnetic field profiles, including spatial non-uniformities, field gradients, and asymmetric distributions caused by nearby ferromagnetic materials, imaging equipment, or electromagnetic interference. The learning-based model’s performance under such irregular field conditions remains untested and would likely degrade without prior exposure to these variations during training. Robust clinical deployment would require expanding the training dataset to encompass diverse magnetic field profiles representative of surgical environments, including both measured clinical field distributions and systematically varied asymmetric patterns. This expanded training approach would enable the model to learn invariant features across field variations rather than overfitting to idealized uniform fields.
5 CONCLUSION
@@
Funding Source Statement
This work was supported by the Canadian Institutes of Health Research (CIHR) Project Grant 452287.
Declaration of Competing Interest
The authors declare that they have no competing financial interests or personal relationships that could have influenced the work reported in this paper.
Declaration of Generative AI Use
The author(s) used ChatGPT for grammar and language checking only. No scientific content was generated using AI. The author(s) reviewed and edited the content and take full responsibility for the published article.
Data Availability
Data will be made available on request.